OpenStudy (yacoub1993):

How do i identify. and where to find theta on the scientific calculator.. Verify the identity. cot θ ∙ sec θ = csc θ

OpenStudy (homeworksucks):

I don't think a calculator has a button for theta..

OpenStudy (yacoub1993):

then how to solve for a question like that

OpenStudy (homeworksucks):

Its more an expression than an equation, you manipulate the left side to get the right one

OpenStudy (yacoub1993):

@Homeworksucks Show me step by step because i m new to this

OpenStudy (mertsj):

\[\frac{\cos \theta}{\sin \theta}\times \frac{1}{\cos \theta}=\] \[\frac{1}{\sin \theta}=\csc \theta \]

OpenStudy (anonymous):

start with equivalent expressions for each... \[\cot \theta = \frac{ \cos \theta }{ \sin \theta }\text{, }\sec \theta = \frac{ 1}{\cos \theta }\text{, and }\csc \theta = \frac{ 1}{\sin \theta }\] substitute in those expressions and see what you get... start on the left and see if you get the right.

OpenStudy (homeworksucks):